OpenStudy (anonymous):

13r+10=5r-46 need a solution

OpenStudy (katrinakaif):

Bring the 5r to the other side by substracting it giving 13r - 5 + 10 = -46 Bring the 10r to the other side by substracting it which is 13r - 5r = 10 - 46 collect like terms 8r = -36 r = -36/8 simplify r = -9/2

OpenStudy (anonymous):

no i don't think so. this is the danger in "bringing" things instead of adding or subtracting.

OpenStudy (anonymous):

\[13r+10=5r-46\]\subtract 5r from both sides \[13r-5r+10=-46\] \[8r+10=-46\] subtract 10 from both sides \[8r=-46-10\] \[8r=-56\] divide by 8 \[r=\frac{-56}{8}\] \[r=-7\]

OpenStudy (anonymous):

just to make it clear, in math we add, subtract, multiply and divide. we do not "bring"